### v2

@@ 1.1 Our results
A function $f:\mathcal{G}\to\{-1,+1\}$ is $\epsilon$-far from $s$-sparse-Boolean if for every $s$-sparse function $g:\mathcal{G}\to\{-1,+1\}$ the $\ell_{2}$-distance of $f$ and $g$ is at least $\sqrt{\epsilon}$. [^5] In property testing usually the distance measure used is Hamming distance between two Boolean functions. But since we are using $\ell_{2}$ distance in our other theorem, so for ease of presentation, we have defined the farness in terms of $\ell_{2}$ instead of Hamming distance. Also, note that for a pair of Boolean-valued functions the square of the $\ell_{2}$ distance and Hamming distance are the same up to a multiplicative factor of 44, see Lemma 2.43.
@@ 1.2 Proof ideas
In this section, we briefly outline the proof ideas of our main theorems. Note that $\mathbb{Z}_{p}^{n}$ is a vector space for prime pp. For general Abelian groups it can be shown that vector space structure does not exists. To deal with this issue we define an operation $\mathcal{G}\times\mathcal{G}\to\mathbb{Z}_{\Delta}$, which is a generalization of the inner product when $\mathcal{G}=\mathbb{Z}_{p}^{n}$. Note, that the operation, ‘*’, cannot be an inner-product in general as $\mathcal{G}$ does not have a vector space structure. We call it the “pseudo inner-product” operation (see Definition 2.12), and using this pseudo inner-product operation we define certain normal subgroups that acts like subspaces of $\mathcal{G}$, and their cosets are used to partition $\mathcal{G}$. Details are given in Section 2. Let us first sketch the proof of Theorem 1.3.
@@
To show the above, we first utilize a probabilistic method to prove that for each character $\chi_{i}\in B$ in the Fourier support of $f$, there exists a matrix $A\in\mathbb{Z}_{\Delta}^{k\times T}$, and a column vector $b\in\mathbb{Z}_{\Delta}^{T\times 1}$, such that – (1) $\chi_{i}$ is a solution of the system of equations $A\chi=b$, and (2) no other character in the Fourier support of $f$ is a solution of $A\chi=b$, where $B$ be the set of $s$-largest Fourier coefficients of $f$. Using our pseudo inner product structure, we establish the existence of such AA and $b$, we consider the Fourier transform of the projection (Definition 3.1) for the solution space of $A\chi=b$ (see Definition 3.1 and Lemma 3.2). The projection operator, as the name suggests, is an operator that projects $\mathcal{G}$ onto a coset which yields a partition of the Fourier spectrum of $f$. Then we show that the $\ell_{2}$ Fourier weight of $S\cap H$, i.e., $\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}$ is upper bounded by $\mu/\sqrt{s}$, where $S=\overline{B}$, and $H$ is a coset of $A^{\perp}$ that are solutions to the system of linear equations $A\chi=b$. For details, see Section 3.
@@
We compute $\mathbb{E}[G(x)^{2}]$ in order to find an upper bound on the Fourier coefficients of H:=G(2fG)H:=G(2f-G), which helps us to conclude that $\widehat{F^{2}}(\chi)=0$ for all $\chi\neq\chi_{0}$, and $\widehat{F^{2}}(\chi_{0})=1$, where $\chi_{0}$ is the character which takes the value 11 at all points in $\mathcal{G}$. Then we complete the proof by showing that the $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$ is $\leq\mu^{2}/2$, which implies that $F$ is $\sqrt{2}\mu$ close to $f$ in $\ell_{2}$ by Lemma 2.43. This idea has also been employed in  [GOS+11].
@@
$V_{b,r_{1},\ldots,r_{k}}\;:=\;\{x\in\mathcal{G}:r_{j}\cdot x\,=\,b_{j}\leavevmode\nobreak\ \pmod{\Delta}\ \forall j\in[t]\}\;.$
@@
where $s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}$ is the Fourier sparsity of the function $f|_{V_{b,r_{1},\ldots,r_{t}}}$. [GOS+11] defined the degree over $\mathbb{Z}_{2}^{n}$ via similar restrictions. However, [GOS+11] argued that this is a natural definition of the degree over $\mathbb{Z}_{2}^{n}$. To argue, observe that $\widehat{f}=\frac{1}{2^{n}}H_{n}f$, where $H_{n}$ is the $2^{n}\times 2^{n}$ Hadamard matrix, when $x_{i}\in\mathbb{Z}_{2}^{n}$ are seen in lexicographic order. Consider the restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ that takes the value 11 at those points such that each entry of $\widehat{f}|_{V_{b,r_{1},\ldots,r_{t}}}$ is nonzero. Then $\deg_{2}$ can be defined as the dimension of $V_{b,r_{1},\ldots,r_{t}}$ which is largest amongst them! In that case, all the Fourier coefficients of $f|_{V_{b,r_{1},\ldots,r_{t}}}$ are nonzero.
@@ 2.2 The subgroup $H^{\perp}$
where $\omega_{N}$ is a primitive $N^{th}$ root of unity of order NN, and $|H|$ denotes the order of the subgroup $H$.
@@ 2.4 Basics from algebraic number theory
Recall that the Galois group, denoted $\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})$ of automorphisms of $Q(\omega_{n})$; it is isomorphic to the multiplicative group $\mathbb{Z}_{n}^{\ast}$, of integers $\bmod\leavevmode\nobreak\ n$, i.e., all integers $k$ such that gcd(k,n)=1(k,n)=1. Note that $|\mathbb{Z}_{n}^{\ast}|=\varphi(n)$, where $\varphi$ is the Euler’s totient function, where $\varphi(n):=n\cdot\prod_{\text{prime}\leavevmode\nobreak\ p\mid n}(1-1/p)$. In simple words, think of $\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})$, as consisting $\varphi(n)$ many automorphism $\sigma_{k}$, that sends $\omega_{n}\mapsto\omega_{n}^{k}$, where $k\in\mathbb{Z}_{n}^{\ast}$. For basics, we refer to [Gre20, Chapter 1].
@@ 2.5 Miscellaneous
So, the square of the $\ell_{2}$ distance between $f$ and $g$ is equal to 44 times $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$. ∎
@@ 3.1 The projection on a coset
Set $k=\lceil\log_{\Delta}2s\rceil$, and let AA be a $k\times t$ matrix with the entries $a_{ij}$ such that $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$, and $b$ be a $k\times 1$ matrix.
Let $E$ be the event that the matrix AA is chosen such that its entries $a_{ij}$ are invertible in $\mathbb{Z}_{\Delta}$ for all $i\geq j$, and $a_{ij}$ are chosen independently and uniformly at random.
@@
We need to show that for each $\chi_{r_{i}}\in B$, there exists a $k\times n$ matrix AA, with the entries $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$, and a $k\times 1$ matrix $b$ such that
@@
To show the existence of AA and $b$, given $\chi_{r_{i}}\in B$, for which Eq. 3.4 is true we will need to consider two cases:
@@
. Observe that for all AA, we have $A\chi_{r_{i}}=0$, so $\Pr_{A}[A\chi_{r_{i}}=0]=1$. If entries of AA are sampled independently and uniformly at random from $\mathbb{Z}_{\Delta}$, with $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$, then, for all $\chi_{r_{j}}\in B\setminus\{\chi_{r_{i}}\}$, we have
@@
Note that $\Pr_{A}\left[E_{1}\right]$ denotes the probability of event $E_{1}$ when AA is the random matrix generated by the above-mentioned procedure. Using union bound, we get
@@
Therefore, if $\chi_{r_{i}}=0$ then there exists a $k\times t$ matrix AA, with the entries $a_{ij}$ are invertible for all $i\geq j$ such that
@@
. Fix a nonzero element $b\in\mathcal{G}\setminus\{(0,\dots,0)\}$. Again, let AA be a random $k\times t$ matrix with entries of AA being sampled independently and uniformly at random from $\mathbb{Z}_{\Delta}$ and the entries $a_{ij}$ are invertible in $\mathbb{Z}_{\Delta}$ for all $i\geq j$. If $\chi_{r_{j}}=\alpha\chi_{r_{i}}$, where $\alpha\in\mathbb{Z}_{\Delta}$ and $\gcd(\alpha-1,\Delta)=1$, then
@@
Note that $A_{\ell}$ and $b_{\ell}$ denote the $\ell$-th row of AA and $\ell$-th coordinate of $b$ respectively, and $[k]$ denotes the set $\{1,\dots,k\}$. Observe that the last equality in the above equation follows from the fact that rows of AA are independent and uniformly random samples of $\mathbb{Z}_{\Delta}^{k}$. Therefore
@@
From the above two cases we can now say that given any $\chi_{r_{i}}\in B$, there exists a $k\times t$ matrix AA, with the entries $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$ and a $k\times 1$ matrix $b$ such that
@@
which is of the form $\frac{g(\omega_{\Delta})}{\Delta^{k}}$, since by Lemma 2.27, $|(A^{\perp})^{\perp}|=\frac{|\mathbb{Z}_{\Delta}^{T}|}{|A^{\perp}|}=\Delta^{k}$ (since we have sampled the entries of the matrix AA from $\mathbb{Z}_{\Delta}$), where $g(X)\in\mathbb{Z}[X]$ is a nonzero polynomial with degree $\leq\Delta-1$. So $P_{H}f(x)$ is $k$-granular.
@@ 3.3 Lower bound on the Fourier coefficients
where $\chi_{0}$ is the character which takes the value 11 at all points in $\mathcal{G}$, as defined in Lemma 2.4, part 1.
@@ 4 A $\Delta$-independent lower bound
Now, we prove a lower bound on the Fourier coefficient which is $\Delta$-independent; this is an immediate consequence of Corollary $4.$
@@
We know that $s_{f}\geq\varphi(\Delta)$, by Corollary $4.$ Hence, we combine this with Theorem 1.5 , to get
@@ 6.1 The Algorithm
We perform a random shift $u\in\mathbb{Z}_{\Delta}^{t}$6.9) and rename the buckets $C(b)$ as $C_{u}(b)$. Observe that the buckets $C_{u}(b)$ gives a partition of $\mathcal{G}$ and each $C_{u}(b)$ corresponds to a coset in $\mathcal{G}/H^{\perp}$. Thus once we have picked the random vector space $H$ in lines 4 and 5 we have a partition of $\mathcal{G}$ (and thus the Dual space of $\mathcal{G}$) into $\Delta^{t}$ buckets where each bucket is of the form $(r+H^{\perp})$ for $r\in\mathcal{G}$.
@@
where $v_{i}$ are the rows of the matrix AA.
@@
Let us define the random variables $Y_{b}=\mathsf{wt}(C_{u}(b)\cap A_{2})$ for each $b$. Then $Y_{b}=\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\cdot I_{\chi\to b+u}$, where $I_{\chi\to b+u}$ is the indicator random variable which takes the value 11 if $\chi\in C_{u}(b)$ and 0 otherwise.
@@ 7 Degree over $\mathcal{G}$
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean valued function. Let us arrange the elements $x\in\mathcal{G}$ in lexicographic order along the rows and columns of the matrix $V_{1}\otimes\cdots\otimes V_{T}$. Consider all possible restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ of $f$, where $V_{b,r_{1},\ldots,r_{t}}$ is a coset of $\mathcal{G}$ as defined in Definition 2.17. Then the degree over $\mathcal{G}$ of $f$ is defined as the dimension of the largest coset $V_{b,r_{1},\ldots,r_{t}}$ such that the restriction $f|_{V_{b,r_{1},\ldots,r_{t}}}$ takes the value 11 at those points such that each entry of the row/column matrix formed by the sum of the rows/columns corresponding to the elements where $f|_{V_{b,r_{1},\ldots,r_{t}}}$ takes the value 11 is $\neq 0$.
@@
Note that, the first row/column, corresponding to the constant coefficient, is not equal to 0. In the case when $\mathcal{G}$ is of prime order, the constant coefficient is always $\neq 0$, as $f$ can take the value 11 at an odd number of elements and the value 1-1 at an even number of elements, or vice versa. In case of general Abelian groups, this may not be true, so the restriction has to be chosen in such a way that it does not take the value 11 at half of the points, so that the constant coefficient is not equal to 0.

### v3

@@ 1.1 Our results
A function $f:\mathcal{G}\to\{-1,+1\}$ is $\epsilon$-far from $s$-sparse-Boolean if for every $s$-sparse function $g:\mathcal{G}\to\{-1,+1\}$ the $\ell_{2}$-distance of $f$ and $g$ is at least $\sqrt{\epsilon}$. [^5] In property testing usually the distance measure used is Hamming distance between two Boolean functions. But since we are using $\ell_{2}$ distance in our other theorem, so for ease of presentation, we have defined the farness in terms of $\ell_{2}$ instead of Hamming distance. Also, note that for a pair of Boolean-valued functions the square of the $\ell_{2}$ distance and Hamming distance are the same up to a multiplicative factor of $4$, see Lemma 2.43.
@@ 1.2 Proof ideas
In this section, we briefly outline the proof ideas of our main theorems. Note that $\mathbb{Z}_{p}^{n}$ is a vector space for prime $p$. For general Abelian groups it can be shown that vector space structure does not exists. To deal with this issue we define an operation $\mathcal{G}\times\mathcal{G}\to\mathbb{Z}_{\Delta}$, which is a generalization of the inner product when $\mathcal{G}=\mathbb{Z}_{p}^{n}$. Note, that the operation, ‘*’, cannot be an inner-product in general as $\mathcal{G}$ does not have a vector space structure. We call it the “pseudo inner-product” operation (see Definition 2.12), and using this pseudo inner-product operation we define certain normal subgroups that acts like subspaces of $\mathcal{G}$, and their cosets are used to partition $\mathcal{G}$. Details are given in Section 2. Let us first sketch the proof of Theorem 1.3.
@@
To show the above, we first utilize a probabilistic method to prove that for each character $\chi_{i}\in B$ in the Fourier support of $f$, there exists a matrix $A\in\mathbb{Z}_{\Delta}^{k\times T}$, and a column vector $b\in\mathbb{Z}_{\Delta}^{T\times 1}$, such that – (1) $\chi_{i}$ is a solution of the system of equations $A\chi=b$, and (2) no other character in the Fourier support of $f$ is a solution of $A\chi=b$, where $B$ be the set of $s$-largest Fourier coefficients of $f$. Using our pseudo inner product structure, we establish the existence of such $A$ and $b$, we consider the Fourier transform of the projection (Definition 3.1) for the solution space of $A\chi=b$ (see Definition 3.1 and Lemma 3.2). The projection operator, as the name suggests, is an operator that projects $\mathcal{G}$ onto a coset which yields a partition of the Fourier spectrum of $f$. Then we show that the $\ell_{2}$ Fourier weight of $S\cap H$, i.e., $\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}$ is upper bounded by $\mu/\sqrt{s}$, where $S=\overline{B}$, and $H$ is a coset of $A^{\perp}$ that are solutions to the system of linear equations $A\chi=b$. For details, see Section 3.
@@
We compute $\mathbb{E}[G(x)^{2}]$ in order to find an upper bound on the Fourier coefficients of H:=G(2fG)H:=G(2f-G), which helps us to conclude that $\widehat{F^{2}}(\chi)=0$ for all $\chi\neq\chi_{0}$, and $\widehat{F^{2}}(\chi_{0})=1$, where $\chi_{0}$ is the character which takes the value $1$ at all points in $\mathcal{G}$. Then we complete the proof by showing that the $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$ is $\leq\mu^{2}/2$, which implies that $F$ is $\sqrt{2}\mu$ close to $f$ in $\ell_{2}$ by Lemma 2.43. This idea has also been employed in [GOS+11].
@@
$V_{b,r_{1},\ldots,r_{k}}\;:=\;\{x\in\mathcal{G}:r_{j}\cdot x\,=\,b_{j}\penalty 10000\ \pmod{\Delta}\ \forall j\in[t]\}\;.$
@@
where $s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}$ is the Fourier sparsity of the function $f|_{V_{b,r_{1},\ldots,r_{t}}}$. [GOS+11] defined the degree over $\mathbb{Z}_{2}^{n}$ via similar restrictions. However, [GOS+11] argued that this is a natural definition of the degree over $\mathbb{Z}_{2}^{n}$. To argue, observe that $\widehat{f}=\frac{1}{2^{n}}H_{n}f$, where $H_{n}$ is the $2^{n}\times 2^{n}$ Hadamard matrix, when $x_{i}\in\mathbb{Z}_{2}^{n}$ are seen in lexicographic order. Consider the restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ that takes the value $1$ at those points such that each entry of $\widehat{f}|_{V_{b,r_{1},\ldots,r_{t}}}$ is nonzero. Then $\deg_{2}$ can be defined as the dimension of $V_{b,r_{1},\ldots,r_{t}}$ which is largest amongst them! In that case, all the Fourier coefficients of $f|_{V_{b,r_{1},\ldots,r_{t}}}$ are nonzero.
@@ 2.2 The subgroup $H^{\perp}$
where $\omega_{N}$ is a primitive $N^{th}$ root of unity of order $N$, and $|H|$ denotes the order of the subgroup $H$.
@@ 2.4 Basics from algebraic number theory
Recall that the Galois group, denoted $\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})$ of automorphisms of $Q(\omega_{n})$; it is isomorphic to the multiplicative group $\mathbb{Z}_{n}^{\ast}$, of integers $\bmod\penalty 10000\ n$, i.e., all integers $k$ such that gcd(k,n)=1(k,n)=1. Note that $|\mathbb{Z}_{n}^{\ast}|=\varphi(n)$, where $\varphi$ is the Euler’s totient function, where $\varphi(n):=n\cdot\prod_{\text{prime}\penalty 10000\ p\mid n}(1-1/p)$. In simple words, think of $\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})$, as consisting $\varphi(n)$ many automorphism $\sigma_{k}$, that sends $\omega_{n}\mapsto\omega_{n}^{k}$, where $k\in\mathbb{Z}_{n}^{\ast}$. For basics, we refer to [Gre20, Chapter 1].
@@ 2.5 Miscellaneous
So, the square of the $\ell_{2}$ distance between $f$ and $g$ is equal to $4$ times $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$. ∎
@@ 3.1 The projection on a coset
Set $k=\lceil\log_{\Delta}2s\rceil$, and let $A$ be a $k\times t$ matrix with the entries $a_{ij}$ such that $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$, and $b$ be a $k\times 1$ matrix.
Let $E$ be the event that the matrix $A$ is chosen such that its entries $a_{ij}$ are invertible in $\mathbb{Z}_{\Delta}$ for all $i\geq j$, and $a_{ij}$ are chosen independently and uniformly at random.
@@
We need to show that for each $\chi_{r_{i}}\in B$, there exists a $k\times n$ matrix $A$, with the entries $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$, and a $k\times 1$ matrix $b$ such that
@@
To show the existence of $A$ and $b$, given $\chi_{r_{i}}\in B$, for which Eq. 3.4 is true we will need to consider two cases:
@@
. Observe that for all $A$, we have $A\chi_{r_{i}}=0$, so $\Pr_{A}[A\chi_{r_{i}}=0]=1$. If entries of $A$ are sampled independently and uniformly at random from $\mathbb{Z}_{\Delta}$, with $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$, then, for all $\chi_{r_{j}}\in B\setminus\{\chi_{r_{i}}\}$, we have
@@
Note that $\Pr_{A}\left[E_{1}\right]$ denotes the probability of event $E_{1}$ when $A$ is the random matrix generated by the above-mentioned procedure. Using union bound, we get
@@
Therefore, if $\chi_{r_{i}}=0$ then there exists a $k\times t$ matrix $A$, with the entries $a_{ij}$ are invertible for all $i\geq j$ such that
@@
. Fix a nonzero element $b\in\mathcal{G}\setminus\{(0,\dots,0)\}$. Again, let $A$ be a random $k\times t$ matrix with entries of $A$ being sampled independently and uniformly at random from $\mathbb{Z}_{\Delta}$ and the entries $a_{ij}$ are invertible in $\mathbb{Z}_{\Delta}$ for all $i\geq j$. If $\chi_{r_{j}}=\alpha\chi_{r_{i}}$, where $\alpha\in\mathbb{Z}_{\Delta}$ and $\gcd(\alpha-1,\Delta)=1$, then
@@
Note that $A_{\ell}$ and $b_{\ell}$ denote the $\ell$-th row of $A$ and $\ell$-th coordinate of $b$ respectively, and $[k]$ denotes the set $\{1,\dots,k\}$. Observe that the last equality in the above equation follows from the fact that rows of $A$ are independent and uniformly random samples of $\mathbb{Z}_{\Delta}^{k}$. Therefore
@@
From the above two cases we can now say that given any $\chi_{r_{i}}\in B$, there exists a $k\times t$ matrix $A$, with the entries $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$ and a $k\times 1$ matrix $b$ such that
@@
which is of the form $\frac{g(\omega_{\Delta})}{\Delta^{k}}$, since by Lemma 2.27, $|(A^{\perp})^{\perp}|=\frac{|\mathbb{Z}_{\Delta}^{T}|}{|A^{\perp}|}=\Delta^{k}$ (since we have sampled the entries of the matrix $A$ from $\mathbb{Z}_{\Delta}$), where $g(X)\in\mathbb{Z}[X]$ is a nonzero polynomial with degree $\leq\Delta-1$. So $P_{H}f(x)$ is $k$-granular.
@@ 3.3 Lower bound on the Fourier coefficients
where $\chi_{0}$ is the character which takes the value $1$ at all points in $\mathcal{G}$, as defined in Lemma 2.4, part 1.
@@ 4 A $\Delta$-independent lower bound
Now, we prove a lower bound on the Fourier coefficient which is $\Delta$-independent; this is an immediate consequence of Corollary 4.4.
@@
We know that $s_{f}\geq\varphi(\Delta)$, by Corollary 4.4. Hence, we combine this with Theorem 1.5 , to get
@@ 6.1 The Algorithm
We perform a random shift $u\in\mathbb{Z}_{\Delta}^{t}$ ( 6.9) and rename the buckets $C(b)$ as $C_{u}(b)$. Observe that the buckets $C_{u}(b)$ gives a partition of $\mathcal{G}$ and each $C_{u}(b)$ corresponds to a coset in $\mathcal{G}/H^{\perp}$. Thus once we have picked the random vector space $H$ in lines 4 and 5 we have a partition of $\mathcal{G}$ (and thus the Dual space of $\mathcal{G}$) into $\Delta^{t}$ buckets where each bucket is of the form $(r+H^{\perp})$ for $r\in\mathcal{G}$.
@@
where $v_{i}$ are the rows of the matrix $A$.
@@
Let us define the random variables $Y_{b}=\mathsf{wt}(C_{u}(b)\cap A_{2})$ for each $b$. Then $Y_{b}=\sum_{\chi\in A_{2}}|\widehat{f}(\chi)|^{2}\cdot I_{\chi\to b+u}$, where $I_{\chi\to b+u}$ is the indicator random variable which takes the value $1$ if $\chi\in C_{u}(b)$ and 0 otherwise.
@@ 7 Degree over $\mathcal{G}$
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean valued function. Let us arrange the elements $x\in\mathcal{G}$ in lexicographic order along the rows and columns of the matrix $V_{1}\otimes\cdots\otimes V_{T}$. Consider all possible restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ of $f$, where $V_{b,r_{1},\ldots,r_{t}}$ is a coset of $\mathcal{G}$ as defined in Definition 2.17. Then the degree over $\mathcal{G}$ of $f$ is defined as the dimension of the largest coset $V_{b,r_{1},\ldots,r_{t}}$ such that the restriction $f|_{V_{b,r_{1},\ldots,r_{t}}}$ takes the value $1$ at those points such that each entry of the row/column matrix formed by the sum of the rows/columns corresponding to the elements where $f|_{V_{b,r_{1},\ldots,r_{t}}}$ takes the value $1$ is $\neq 0$.
@@
Note that, the first row/column, corresponding to the constant coefficient, is not equal to 0. In the case when $\mathcal{G}$ is of prime order, the constant coefficient is always $\neq 0$, as $f$ can take the value $1$ at an odd number of elements and the value 1-1 at an even number of elements, or vice versa. In case of general Abelian groups, this may not be true, so the restriction has to be chosen in such a way that it does not take the value $1$ at half of the points, so that the constant coefficient is not equal to 0.